OpenStudy (anonymous):

Find the exact circumference of a circle with an area equal to 36 sq. in.

OpenStudy (nurali):

A = pi r^2 sqrt(A/pi) = r C = 2r pi C = 2sqrt(A/pi) * pi

OpenStudy (anonymous):

is the answer 12 times pie then

OpenStudy (anonymous):

alright thanks a lot

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Area=36 pi*r*r=36 r*r=11.14 r=3.38 C=2piR 2*3.14*3.38=21.26sq in.